OpenStudy (anonymous):

For f(x)= 5x + 1, find f(–4).

OpenStudy (mathstudent55):

Just replace x with -4 and calculate.

OpenStudy (mathstudent55):

f(x) = 5x + 1 f(-4) = 5(-4) + 1 = ...

OpenStudy (anonymous):

simply replace the x in the problem with \(-4\) basically, the question asks that if the function of \(x\) is \(5x+1\) then what is the function of \(-4\) so you would do as such:\[f(x)=5x+1\]\[f(-4)=5(-4)+1\]\[f(-4)=-20+1\]\[f(-4)=-19\]